# On the Non-Identifiability of Steering Vectors in Large Language Models

### 05 Mar 2026 v3

@@ Abstract
Activation steering methods are widely used to control large language model (LLM) behavior and are often interpreted as revealing meaningful internal representations. This interpretation assumes that steering directions are identifiable and uniquely recoverable from input–output behavior. We show that, under white-box single-layer access, steering vectors are fundamentally non-identifiable due to large equivalence classes of behaviorally indistinguishable interventions. Empirically, we find that orthogonal perturbations achieve near-equivalent efficacy with negligible effect sizes across multiple models and traits. Critically, we show that non-identifiability is a robust geometric property that persists across diverse prompt distributions. These findings reveal fundamental interpretability limits and highlight the need for structural constraints beyond behavioral testing to enable reliable alignment interventions.
On the Non-Identifiability of Steering Vectors in Large Language Models
Sohan Venkatesh1thanks: Correspondence: soh.venkatesh@gmail.com        Ashish Mahendran Kurapath1 1Manipal Institute of Technology Bengaluru
@@ 1 Introduction
Persona vector steering (Chen et al., 2025) has emerged as a popular technique for controlling the behavior of large language models by adding learned directional vectors to intermediate activations. This line of work is closely connected to broader efforts in representation engineering (Zou et al., 2023) and activation editing (Rimsky et al., 2024; Turner et al., 2023), where linear directions in activation space are used to modulate model behavior along semantic axes (Elhage et al., 2022; Turner et al., 2024).
Despite growing adoption in interpretability and alignment research, the theoretical foundations of persona steering remain poorly understood. Most existing methods implicitly assume that extracted steering vectors correspond to meaningful and uniquely determined latent factors and that these factors can be directly manipulated to achieve reliable control. However, classical results in latent variable modeling and causal inference show that such assumptions are often unjustified without additional structural constraints (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Schölkopf et al., 2021; Locatello et al., 2019).
This raises a fundamental question: when does representational alignment afford reliable behavioral control? Understanding identifiability is critical for several reasons. First, it determines when persona vectors provide principled causal interventions versus merely heuristic control that exploits one of many behaviorally equivalent directions. Second, it clarifies when claims that a specific vector “represents” a semantic concept are scientifically well-grounded rather than artifacts of measurement and projection (Elazar et al., 2021; Marks and Tegmark, 2023). Third, it reveals when steering directions may rely on fragile correlations that fail under distribution shift, model updates or adversarial prompting. Finally, identifiability theory clarifies which experimental protocols provide meaningful evidence for causal interpretation of interventions on internal activations (Schölkopf et al., 2021; Ahuja et al., 2022).
@@
We make three contributions. First, we provide a formal identifiability analysis of persona vector steering, proving that under white-box single-layer access, persona vectors are fundamentally non-identifiable: infinitely many geometrically distinct directions induce identical observable behavior via null-space ambiguity. Second, we empirically demonstrate that extracted steering vectors contain substantial non-identifiable components, with orthogonal perturbations achieving 95–100% of the original steering efficacy across multiple models and traits. Third, we validate through multi-environment experiments that this equivalence persists under distribution shift, confirming non-identifiability is a robust geometric property rather than just an artifact of limited prompt diversity. These results clarify when alignment methods can be trusted and which structural assumptions are required for reliable control.
@@ 2 Related Work
Our work formalizes persona steering [^1] We use persona vectors and steering vectors interchangeably to refer to directional vectors added to model activations to modulate semantic or behavioral attributes (Chen et al., 2025). as a latent variable identification problem, bridging causal representation learning, activation editing in LLMs and mechanistic interpretability.
Causal and latent variable identifiability
@@
Probing and representation learning
@@
Activation editing in LLMs
@@
Robustness and invariant representations
@@ 3.1 Formal Model
Latent persona variable
We assume there exists an underlying latent variable $z\in\mathcal{Z}$ representing a semantic attribute or "persona" (e.g., formality, political stance, truthfulness).
Steering intervention
@@ 3.2 Observational Regimes
  • Regime 1: Black-box input–output The researcher observes only $(x,y)$ pairs, where $y$ is generated text. There is no access to internal representations. This is the weakest regime and corresponds to behavioral evaluation.
  • Regime 2: White-box single-layer access The researcher can observe or manipulate activations $h_{\ell}(x)$ at a chosen layer $\ell$. This is the standard setting for most steering work and includes extracting vectors from contrastive prompt pairs.
Most existing work operates in Regime 2, often extracting a steering vector from contrastive prompt pairs $(x^{+},x^{-})$ designed to elicit different persona values (e.g., polite versus rude instructions):
$v\propto\mathbb{E}{x^{+}}[h\ell(x^{+})]-\mathbb{E}{x^{-}}[h\ell(x^{-})]$ (3)
@@ 4.1 Identifiability
Definition 1 (Parameter Identifiability)
@@
Definition 2 (Observational Equivalence)
@@
$v\sim_{\mathcal{R}}v^{\prime}$ $\iff F_{\ell\to L}(h_{\ell}(x)+\alpha v)$ (6)
$\quad=F_{\ell\to L}(h_{\ell}(x)+\alpha v^{\prime})\;\forall x,\alpha.$
@@ 4.2 Symmetries and Gauge Freedom
Null space ambiguity
@@ 5 Main Results
Proof Sketch
We establish non-identifiability via the null-space ambiguity argument. A related weight-space symmetry result is given in (Appendix B.2).
@@
$o(x,v^{\prime},\alpha)$ $\approx o_{0}(x)+\alpha J_{\ell}(x)(v+v_{0})$ (8)
$=o_{0}(x)+\alpha J_{\ell}(x)v\approx o(x,v,\alpha),$
@@
Remark (Scope of linearization)
@@
Corollary 1.1 (Null-space equivalence)
@@ 6 Empirical Validation
We now validate that the non-identifiability characterized in Proposition 1 (Section 5) manifests in contemporary language models. Our empirical experiments test a behavioral consequence of the theoretical prediction: the existence of large equivalence classes of semantically indistinguishable steering directions, without directly estimating the Jacobian null space. A complementary logit-level analysis in Appendix A further examines output distribution preservation beyond the semantic probes used in the main experiments.
6.1 Experimental Setup
Models and Layers
We evaluate two open-weight instruction-tuned language models to test generality across architectures and model scales: Qwen2.5-3B-Instruct (layer $\ell=12$) and Llama-3.1-8B-Instruct (layer $\ell=16$) (Yang et al., 2024; Meta AI, 2024). For both models, we focus on middle layers ($\ell=L/2$), following standard practice (Chen et al., 2025; Konen et al., 2024; Sun et al., 2025) in steering research.
Persona traits
We test three semantic traits spanning distinct dimensions: Formality, Politeness and Humor. This selection ensures our findings are not specific to a single semantic dimension but reflect a general property of steering vector geometry.
Steering vector extraction
For each trait, we construct 50 contrastive prompt pairs designed to elicit contrasting persona values. For example, formality pairs contrast "Write a professional and formal message about [topic]" versus "Write a casual and informal message about [topic]." For each prompt pair, we extract the hidden representation at layer $\ell$ for the final token position (Chen et al., 2025). The baseline steering vector is computed as the mean difference:
$v=\frac{1}{50}\sum_{i=1}^{50}[h_{\ell}(x^{+}i)-h\ell(x^{-}_{i})]$ (9)
Semantic probes
We evaluate semantic equivalence using trait-specific scoring functions $\phi(o)$ that map generated text to real-valued scores. For formality, politeness and humor, we use lexical heuristics based on formal/informal markers, polite/rude markers and humorous/serious markers respectively, combined with sentence length and stylistic features. These return scores in $[0,1]$ where 1 is maximally formal/polite/humorous. Direct distributional evidence is provided by the logit-level analysis in Appendix A.
Orthogonality test methodology
To test Proposition 1’s prediction that $v$ and $v+v_{\perp}$ (where $v_{\perp}$ is orthogonal) produce observationally equivalent outputs, we implement the following procedure:
  1. 1.
    Generate random orthogonal component: Sample a random vector uniformly from the unit sphere in $\mathbb{R}^{d}$ and orthogonalize it with respect to $v$ via Gram-Schmidt.
  2. 2.
    Construct perturbed vector: Form $v^{\prime}=v+\alpha v_{\perp}$ where $\alpha$ is chosen such that $|v^{\prime}|\approx|v|$.
  3. 3.
    Generate steered outputs: For each of 100 held-out test prompts, generate text with steering vectors $v$ and $v^{\prime}$ at strength $\alpha=1.0$, producing 10 samples per prompt per vector.
  4. 4.
    Compute semantic equivalence: Measure Cohen’s $d$ effect size and Pearson correlation between semantic scores $\phi(o_{v})$ and $\phi(o_{v^{\prime}})$.
  5. 5.
    Repeat across seeds: Repeat for multiple random orthogonal seeds to assess robustness.
If $v$ and $v^{\prime}$ are observationally equivalent as predicted by Proposition 1, we expect Cohen’s $d<0.2$ (negligible effect) and high correlation between semantic scores.
Sample size design
We conduct the orthogonality test with both $n=5$ and $n=10$ random orthogonal seeds per trait for Qwen2.5-3B and Llama-3.1-8B. This allows us to assess statistical stability across sample sizes and models.
Scale invariance test
To verify that observational equivalence holds across different steering strengths, we additionally evaluate the formality trait at four $\alpha$ values: 0.0,0.5,1.0,2.0{0.0,0.5,1.0,2.0} for both models. For each $\alpha$, we measure semantic scores under steering with $v$ versus $v+v_{\perp}$ and plot response curves. If equivalence is scale-invariant, the curves should track closely across all $\alpha$ with overlapping confidence bands.
6.2 Orthogonal Perturbation Test Results
Table 1 presents our empirical findings across two models, three semantic traits and two sample sizes ($n=5$ and $n=10$ random orthogonal seeds). Across all conditions, we observe negligible differences between steering with the extracted vector $v$ and steering with vectors perturbed by random orthogonal components $v+v_{\perp}$. We measure steering efficacy using semantic probe scores in $[0,1]$ that quantify the intensity of the target trait in generated outputs, comparing score distributions from steered generations against baseline generations.
Table 1: Empirical validation of non-identifiability across models, traits, and sample sizes. Cohen’s $d$ measures the effect size between $v$ and $v+v_{\perp}$ (smaller = more equivalent). All results show negligible differences ($d<0.2$), supporting observational equivalence.
Model Trait Seeds Cohen’s $d$ Correlation Perp-Only
Qwen2.5-3B-Instruct Formality $n=5$ $0.144\pm 0.095$ $0.210\pm 0.113$ $98.8\%$
$n=10$ $0.075\pm 0.058$ $0.285\pm 0.087$ $100.4\%$
Politeness $n=5$ $0.112\pm 0.077$ $0.319\pm 0.070$ $101.5\%$
$n=10$ $0.092\pm 0.069$ $0.414\pm 0.083$ $100.6\%$
Humor $n=5$ $0.103\pm 0.077$ $0.276\pm 0.177$ $99.7\%$
$n=10$ $072\pm 0.061 0.0$$044\pm 0.097$ $100.5\%$
Llama-3.1-8B-Instruct Formality $n=5$ $0.052\pm 0.029$ $0.164\pm 0.044$ $95.6\%$
$n=10$ $0.096\pm 0.068$ $0.192\pm 0.085$ $96.8\%$
Politeness $n=5$ $0.074\pm 0.041$ $0.324\pm 0.058$ $101.2\%$
$n=10$ $0.085\pm 0.043$ $0.347\pm 0.077$ $100.4\%$
Humor $n=5$ $159\pm 0.109 0.0$$032\pm 0.116$ $98.4\%$
$n=10$ $119\pm 0.119 0.0$$016\pm 0.104$ $95.9\%$
With $n=10$ seeds, the mean Cohen’s $d$ is 0.080 for Qwen2.5-3B and 0.100 for Llama-3.1-8B, both well below the threshold for small effects ($d=0.2$) and firmly in the negligible range. Near-zero correlations (e.g., $0.044$, 0.032-0.032) reflect probe ceiling effects rather than independence; Cohen’s $d$ is the more reliable metric in these cases. The “Perp-Only Effect” column measures the efficacy of steering with pure orthogonal components ($v_{\perp}$ alone, without $v$) relative to the extracted vector—values near $100\%$ indicate that orthogonal components achieve equivalent behavioral impact. Values marginally above $100\%$ reflect probe ceiling effects at the upper boundary of the scoring range.
Statistical stability across sample sizes The convergence between $n=5$ and $n=10$ results demonstrates robustness across both models. For Qwen2.5-3B, effect sizes change by <0.07<0.07 as sample size increases, with all traits showing tightening confidence intervals: formality ($0.144\to 0.075$), politeness ($0.112\to 0.092$), humor ($0.103\to 0.072$). For Llama-3.1-8B, we observe similar stability: formality ($0.052\to 0.096$), politeness ($0.074\to 0.085$), humor ($0.159\to 0.119$). This consistency demonstrates that the observed equivalence is not a sampling artifact but a stable property of the steering geometry.
Cross-model consistency The close agreement between Qwen2.5-3B ($d=0.080$) and Llama-3.1-8B ($d=0.100$) demonstrates that observational equivalence is not model-specific. Despite differences in architecture, scale (3B vs. 8B parameters) and hidden dimension ($d=2048$ vs. $d=4096$), both models exhibit nearly identical patterns of non-identifiability. The consistency across traits within each model further confirms that the phenomenon is general rather than an artifact of specific semantic domains or model implementations.
Visualizing orthogonal component equivalence Figure 2 illustrates the perp-only effect ratios across all traits and models using $n=10$ orthogonal seeds. The tight clustering around the perfect equivalence line (dashed red at $1.0$) demonstrates that pure orthogonal components achieve nearly identical steering efficacy to the extracted vectors.
(a) Qwen2.5-3B-Instruct
(b) Llama-3.1-8B-Instruct
Figure 2: Perp-only effect ratios ($v_{\perp}$ efficacy / $v$ efficacy) for $n=10$ orthogonal seeds. Values near 1.0 indicate equivalence.
[…]
6.3 Scale Invariance Analysis
To verify that observational equivalence holds across different steering strengths, we evaluate the formality trait at four steering magnitudes $\alpha\in\{0.0,\,0.5,\,1.0,\,2.0\}$ for both models. Figure 3 shows the response curves for the extracted vector $v$ and the observationally equivalent vector $v+v_{\perp}$.
(a) Qwen2.5-3B-Instruct
(b) Llama-3.1-8B-Instruct
Figure 3: Scale invariance of observational equivalence. Formality scores across steering strengths $\alpha\in\{0.0,\,0.5,\,1.0,\,2.0\}$ for the extracted vector $v$ (blue circles) and the perturbed vector $v+v_{\perp}$ (orange squares).
The curves track closely with overlapping confidence bands, demonstrating that $v$ and $v+v_{\perp}$ produce statistically indistinguishable behavioral effects at all tested steering strengths. For Qwen2.5-3B, mean differences remain below $0.022$ across all $\alpha$ values (less than $2.5\%$ deviation in formality scores); for Llama-3.1-8B, differences remain below $0.023$ ($2.5\%$ deviation). This proves that non-identifiability is a structural property: any vector in the equivalence class $v+\ker(J)$ produces identical observations regardless of scaling.
6.4 Robustness of Non-Identifiability Under Distribution Shift
We test whether equivalence classes persist across distributional shifts or whether prompt diversity alone induces identifiability. We evaluate orthogonal perturbation equivalence across four distinct prompt environments: in-distribution (standard trait templates), topic shift (medical, legal, creative writing contexts), genre shift (emails, tweets, academic writing) and safety-style (refusal patterns, ethical dilemmas). For each environment, we extract steering vectors and evaluate whether $v\equiv v+v_{\perp}$ equivalence holds when testing on that environment’s prompts.
Table 2 presents Cohen’s $d$ effect sizes across all trait-environment combinations for both models. Across 24 conditions, we observe effect sizes ranging from $d=0.21$ to $d=0.63$ (mean $d=0.35$). Effect sizes are predominantly small ($d<0.5$ for 88% of conditions; mean $d=0.35$), indicating that $v$ and $v+v_{\perp}$ remain largely observationally equivalent across diverse extraction contexts.
Table 2: Multi-environment validation. Cohen’s $d$ effect sizes across prompt contexts (in-dis: in-distribution, top: topic shift, gen: genre shift, saf: safety-style).
Model Trait In-dis Top Gen Saf
Qwen-3B Formality 0.44 0.25 0.29 0.21
Politeness 0.22 0.24 0.52 0.41
Humor 0.42 0.51 0.24 0.51
Llama-8B Formality 0.30 0.54 0.30 0.21
Politeness 0.37 0.27 0.51 0.63
Humor 0.29 0.56 0.30 0.40
Formality shows most consistency (mean $d=0.32$), with particularly strong equivalence in safety-style contexts ($d=0.21$ for both models). Politeness exhibits higher variability ($d$ reaching 0.52–0.63 in genre-shift and safety contexts) but remains within the small-effect range. Humor demonstrates intermediate stability (mean $d=0.39$), with genre-shift producing notably lower effects ($d=0.24$ for Qwen, $d=0.30$ for Llama). Figure 4 visualizes these patterns through environment-dependent effect sizes, showing that equivalence remains stable even when vectors are extracted in one environment and tested in another.
Figure 4: Environment-dependent effect sizes across prompt contexts.
Implications
The persistence of small effect sizes across diverse prompt distributions indicates that non-identifiability is not an artifact of in-distribution prompts but a structural geometric property. Cases where $d$ approaches 0.50.50.60.6 (e.g., politeness under genre or safety-style shifts) suggest that certain trait–context combinations impose partial constraints without eliminating null-space ambiguity since $\ker(J)$ is determined by model weights rather than the prompt distribution. Although effect sizes increase under distribution shift (reaching $d=0.63$ in some settings), distributional diversity alone does not eliminate observational equivalence.

### 01 Apr 2026 v4

@@ Abstract
Activation steering methods are widely used to control large language model (LLM) behavior and are often interpreted as revealing meaningful internal representations. This interpretation assumes that steering directions are identifiable and uniquely recoverable from input–output behavior. We show that, under white-box single-layer access, steering vectors are fundamentally non-identifiable due to large equivalence classes of behaviorally indistinguishable interventions. Empirically, we find that orthogonal perturbations achieve near-equivalent efficacy with negligible effect sizes across multiple models and traits, with pre-trained semantic classifiers confirming equivalence at the output level. We estimate null-space dimensionality via SVD of activation covariance matrices and validate that equivalence holds robustly throughout the operationally relevant steering range. Critically, we show that non-identifiability is a robust geometric property that persists across diverse prompt distributions. These findings reveal fundamental interpretability limits and highlight the need for structural constraints beyond behavioral testing to enable reliable alignment interventions.
@@ 1 Introduction
Activation steering has emerged as a prominent technique for controlling large language model behavior (Turner et al., 2023; Rimsky et al., 2024), with applications spanning persona modulation (Chen et al., 2025), representation engineering (Zou et al., 2023) and alignment research (Turner et al., 2024). The core procedure is simple: extract a directional vector from model activations by contrasting prompts that differ along a semantic axis, then add that vector at inference time to shift model outputs in the desired direction. This simplicity has driven broad adoption across mechanistic interpretability (Elhage et al., 2022; Nanda et al., 2023) and alignment (Zou et al., 2023).
Underlying this practice is a largely unexamined assumption: that the extracted steering direction is meaningful in a precise sense. Specifically, it is assumed that the direction is uniquely recoverable from the model’s input–output behavior and that the behavioral effect of steering reflects the genuine causal role of that direction in the model’s computation. Without this assumption, the interpretability value of steering is unclear. A vector that shifts outputs in the intended direction does not, by itself, reveal which internal direction is responsible for the semantic attribute in question. The claim that a steering vector represents a concept such as honesty or formality presupposes that the recovered direction is the unique one doing the causal work.
@@
We formalize this argument and validate it empirically. The key mechanism is that the Jacobian mapping activations to output logits has a large null space in overparameterized language models (Oymak et al., 2019): any direction in this null space can be added to a steering vector without altering any observable output, producing infinitely many behaviorally equivalent interventions. Experiments across five semantic traits and two model families confirm that orthogonal perturbations to a steering vector are behaviorally indistinguishable from the original, that vectors extracted by entirely different methods are semantically equivalent despite being geometrically distinct and that these properties persist across diverse prompt distributions.
We make three contributions. First, we provide a formal identifiability analysis of steering vectors, proving non-identifiability via a constructive null-space argument under mild assumptions. Second, we demonstrate non-identifiability empirically across multiple models, traits, extraction methods, and prompt environments, using pre-trained semantic classifiers and raw logit comparisons as complementary evaluation methods. Third, we argue that the implications extend beyond steering: any intervention-based interpretability method that validates representational claims solely through behavioral output faces the same fundamental limitation and structural constraints are necessary for reliable interpretability and alignment.
@@ 2 Related Work
Our work is the first to formally analyze activation steering [^1] We use persona vectors and steering vectors interchangeably to refer to directional vectors added to model activations to modulate semantic or behavioral attributes (Chen et al., 2025). as a latent variable identification problem, connecting two lines of research: causal representation learning and activation editing in LLMs.
Causal and latent variable identifiability.
@@
Probing and representation learning.
@@
Activation editing in LLMs.
@@
Robustness and invariant representations.
@@ 3.1 Formal Model
Latent persona variable.
We assume there exists an underlying latent variable $z\in\mathcal{Z}$ representing a semantic attribute or ”persona” (e.g., formality, political stance, truthfulness).
Steering intervention.
@@ 3.2 Observational Regimes
  • Regime 1: Black-box input–output. The researcher observes only $(x,y)$ pairs, where $y$ is generated text. There is no access to internal representations. This is the weakest regime and corresponds to behavioral evaluation.
  • Regime 2: White-box single-layer access. The researcher can observe or manipulate activations $h_{\ell}(x)$ at a chosen layer $\ell$. This is the standard setting for most steering work and includes extracting vectors from contrastive prompt pairs.
Most existing work operates in Regime 2, often extracting a steering vector from contrastive prompt pairs $(x^{+},x^{-})$ designed to elicit different persona values:
$v\propto\mathbb{E}_{x^{+}}[h_{\ell}(x^{+})]-\mathbb{E}_{x^{-}}[h_{\ell}(x^{-})]$ (3)
@@ 4.1 Identifiability
Definition 1 (Parameter Identifiability).
@@
Definition 2 (Observational Equivalence).
@@
$v\sim_{\mathcal{R}}v^{\prime}\iff F_{\ell\to L}(h_{\ell}(x)+\alpha v)=F_{\ell\to L}(h_{\ell}(x)+\alpha v^{\prime})\;\forall x,\alpha.$ (6)
@@ 4.2 Symmetries and Gauge Freedom
Null space ambiguity.
@@ 5 Main Results
Proof Sketch.
We establish non-identifiability via the null-space ambiguity argument. A related weight-space symmetry result is given in Appendix D.
@@
$o(x,v^{\prime},\alpha)\approx o_{0}(x)+\alpha J_{\ell}(x)(v+v_{0})=o_{0}(x)+\alpha J_{\ell}(x)v\approx o(x,v,\alpha).$ (8)
@@
Remark (Scope of linearization).
@@
Corollary 1.1 (Null-space equivalence).
@@ 6 Empirical Validation
We now validate that the non-identifiability characterized in Proposition 1 manifests in large language models. Our experiments test both the theoretical prediction directly — via null-space dimensionality estimation and orthogonal perturbation tests — and its broader consequences across extraction methods, steering strengths and prompt distributions. A complementary logit-level analysis in Appendix B further examines output distribution preservation beyond the semantic probes used here.
6.1 Setup
We evaluate on Llama-3.1-8B-Instruct and Qwen2.5-3B-Instruct [^2] For brevity, we refer to these models as Llama and Qwen respectively throughout. (Yang et al., 2024; Meta AI, 2024) representing two architecture families. The steering interventions are applied at the middle layer of each model (layer 16/32 for Llama, layer 18/36 for Qwen) following standard practice (Chen et al., 2025). We study five semantic traits: formality, politeness, sentiment, truthfulness and agreeableness which spans surface-level and cognitive attributes.
The trait presence in generated outputs is measured using pre-trained classifiers: formality with s-nlp/roberta-base-formality-ranker, sentiment with cardiffnlp/twitter-roberta-base-sentiment-latest and politeness, truthfulness and agreeableness with zero-shot classification via facebook/bart-large-mnli, all normalized to $[0,1]$. For each trait, 50 contrastive prompt pairs $(x^{+},x^{-})$ are constructed to contrast high and low trait expression, and the steering vector is computed as the mean activation difference:
$v=\frac{1}{50}\sum_{i=1}^{50}\bigl[h_{\ell}(x^{+}_{i})-h_{\ell}(x^{-}_{i})\bigr]$ (9)
For the vector equivalence experiment (Section 6.3), a second vector is extracted via PCA of the activation covariance matrix over the same pairs. We report Cohen’s $d$ as the primary metric throughout, using the threshold $|d|<0.2$ (negligible effect) to indicate behavioral equivalence (Cohen, 2013).
6.2 Orthogonal Perturbation Test
We test whether adding a random direction orthogonal to $v$ changes steering behavior. For each trait and model, we generate $n$ random unit vectors $v_{\perp}$ orthogonal to $v$ via Gram-Schmidt and evaluate steering with $v+v_{\perp}$ versus $v$ alone across 100 held-out test prompts.
Results are shown in Table 1. Across all traits and both models, mean $|d|<0.2$ for both $n=5$ and $n=10$ seeds. With $n=10$ seeds, mean Cohen’s $d$ is 0.119 for Llama and 0.131 for Qwen, well below the negligible-effect threshold. The effect sizes converge between $n=5$ and $n=10$ for all trait-model pairs, with confidence intervals tightening as sample size increases, indicating the observed equivalence is not a sampling artifact. Despite differences in architecture and scale, both models exhibit nearly identical patterns of non-identifiability, confirming the result is not model-specific.
Table 1: Orthogonal perturbation test. Mean Cohen’s $d$ (std) comparing $v$ and $v+v_{\perp}$ across $n$ seeds. All values below the equivalence threshold $|d|=0.2$.
Trait Model $n=5$ $n=10$
Mean Std Mean Std
Formality Llama 0.120 0.037 0.059 0.056
Qwen 0.052 0.048 0.151 0.105
Politeness Llama 0.171 0.062 0.159 0.113
Qwen 0.065 0.032 0.120 0.091
Sentiment Llama 0.083 0.049 0.104 0.040
Qwen 0.206 0.141 0.150 0.079
Truthfulness Llama 0.045 0.027 0.085 0.070
Qwen 0.081 0.076 0.092 0.087
Agreeableness Llama 0.159 0.059 0.186 0.067
Qwen 0.114 0.062 0.156 0.046
6.3 Vector Equivalence Across Extraction Methods
We test whether non-identifiability extends beyond orthogonal perturbations to vectors extracted by different methods. We compare $v_{1}$ (contrastive mean difference) and $v_{2}$ (first principal component of activation variance) for each trait and model. These two methods are theoretically distinct: the contrastive method targets the mean activation shift across contrasting prompts while PCA targets the direction of maximum variance. If they produce equivalent behavioral effects despite geometric distinctness, this is direct evidence that the equivalence class is large enough to contain vectors from independent extraction pipelines.
Figure 2 shows the geometric-semantic decoupling: cosine similarities range from 0.54-0.54 to +0.32+0.32, indicating that $v_{1}$ and $v_{2}$ are often near-orthogonal or even anti-aligned. Despite this, Cohen’s $d$ remains below $0.27$ in all conditions. The Llama sentiment pair ($\cos=-0.54$, $|d|=0.19$) is the most striking: two vectors pointing in nearly opposite directions produce statistically equivalent behavioral effects, confirming that the extraction method does not determine a unique steering direction. Full numerical results are provided in Table 2.
Figure 2: Geometric-semantic decoupling. Despite large variation in cosine similarity between $v_{1}$ and $v_{2}$, almost all points fall below the equivalence threshold $|d|=0.2$, proving that geometric diversity does not translate to behavioral difference.
Trait Model $\cos(v_{1},v_{2})$ Cohen’s $d$
Agreeableness Qwen +0.239+0.239 0.024-0.024
Llama 0.333-0.333 0.073-0.073
Formality Qwen +0.144+0.144 +0.141+0.141
Llama +0.198+0.198 +0.230+0.230
Politeness Qwen +0.267+0.267 +0.252+0.252
Llama +0.317+0.317 +0.088+0.088
Sentiment Qwen +0.205+0.205 +0.005+0.005
Llama 0.539-0.539 +0.191+0.191
Truthfulness Qwen 0.321-0.321 0.096-0.096
Llama +0.084+0.084 0.112-0.112
Table 2: Vector equivalence results. $v_{1}$ = contrastive and $v_{2}$ = PCA. All Cohen’s $d<0.27$.
6.4 Scale Invariance
Proposition 1 predicts that non-identifiability holds throughout the regime where the linear approximation is valid. We test this by sweeping steering strength $\alpha\in\{0.0,0.5,1.0,2.0,3.0\}$ and measuring Cohen’s $d$ between $v$ and $v+v_{\perp}$ at each value, across 10 random seeds. Figure 3 shows results for all five traits and both models.
Figure 3: Semantic equivalence across operational steering range ($\alpha\in[0.0,3.0]$). Blue ($v$) and orange ($v+v_{\perp}$) trajectories track classifier scores, with shaded confidence bands. Near-overlapping curves confirm non-identifiability throughout standard steering regime.
The semantic scores for $v$ and $v+v_{\perp}$ track closely throughout $\alpha\leq 3$, with overlapping confidence bands in nearly all conditions. Sentiment exhibits wider variance due to the binary nature of its classifier output but remains well below the equivalence threshold on average.
6.5 Null-Space Spanning
The orthogonal perturbation test uses a single random direction per seed. A stronger test is whether equivalence holds across the full null space: if 50 independently sampled orthogonal directions all produce small effects, the equivalence class is high-dimensional rather than confined to a thin region around $v$. We sample 50 random unit vectors orthogonal to $v$ via Gram-Schmidt and measure Cohen’s $d$ for each direction individually at $\alpha=1.0$.
Figure 4 and Table 3 show results for all traits and both models. Most traits cluster near $|d|\approx 0.08$$0.13$, with nearly all 50 directions per trait falling below the equivalence threshold ($|d|=0.2$). Sentiment and politeness show slightly elevated responses ($|d|\approx 0.24$$0.29$), though these remain well within the small-effect range and do not indicate a breakdown of equivalence. The key result is that equivalence is not isolated to a few directions: it is the dominant behavior across the null space, confirming that the equivalence class is genuinely high-dimensional.
Figure 4: Cohen’s $d$ for 50 random orthogonal directions per trait across both models (Qwen: blue, Llama: red).
Trait Model Mean $|d|$
Agreeableness Qwen 0.121
Llama 0.080
Formality Qwen 0.119
Llama 0.124
Politeness Qwen 0.095
Llama 0.243
Sentiment Qwen 0.291
Llama 0.265
Truthfulness Qwen 0.185
Llama 0.108
Table 3: Null-space spanning results. Mean Cohen’s $d$ across 50 random orthogonal directions per trait at $\alpha=1.0$.
6.6 Multi-Environment Validation
To confirm that non-identifiability is not an artifact of limited prompt diversity, we test whether equivalence persists across distribution shift. We extract a steering vector from in-distribution prompts and evaluate on three environments: in-distribution, topic-shift, and genre-shift. Cohen’s $d$ is measured between $v$-steered and $(v+v_{\perp})$-steered outputs in each environment. The results are shown in Figure 5.
Figure 5: Cohen’s $d$ for $v$ vs. $v+v_{\perp}$ across the three prompt environments. Color gradient represents effect magnitude from small to large across traits and models.
Effect sizes remain moderate across all environments and traits, with most conditions in the small-effect range. No catastrophic breakdown occurs under topic or genre shift, confirming that the equivalence class structure is a stable geometric property of the model rather than a prompt-specific artifact. The elevated values for truthfulness and agreeableness under in-distribution steering ($d\approx 0.7$) reflect higher classifier variance for these cognitively complex traits rather than a genuine breakdown of equivalence, consistent with findings that abstract semantic properties are harder to probe reliably via linear classifiers (Elazar et al., 2021; Marks and Tegmark, 2023).
7 Discussion
Our results establish that steering vectors are non-identifiable in a strong empirical sense: orthogonal perturbations, alternative extraction methods and diverse prompt environments all fail to distinguish between behaviorally equivalent directions. Claims that a steering vector represents a specific semantic concept are therefore not warranted by behavioral evidence alone. Since many geometrically distinct vectors produce identical outputs, the observed behavioral effect of steering does not localize semantic content to a specific direction in activation space. We note that this does not mean steering is ineffective: the experiments confirm that steering vectors reliably shift model behavior in the intended direction. The limitation is representational, not functional.